# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

### v2

@@ 7.3 Extra Delay $\Delta d$
Similar to how we measured the extra delay $\Delta d$ introduced by GPT-based SelfDefend in §5.3, we now report the average delay $\Delta d$ introduced by the tuned models on Llama-2-7b-chat in Figure 6. Compared to earlier results in Figure 3, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.008 seconds. In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.072 seconds (0.026 seconds) under AlpacaEval. This is likely because the parameters of fine-tuned models are less than those of GPT-3.5/4. Besides the normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under various jailbreak prompts. For example, the maximum $\Delta d$ now decreases from the earlier 1.56 seconds on GPT-4 to 0.27 seconds. Indeed, except for DAN, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend. These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for potential deployment.
@@ 7.4 Explainability
Figure 5 demonstrate the efficacy of the tuned models in identifying harmful content within jailbreak prompts. Here, we further quantitatively assess whether the identified harmful content actually aligns with the original jailbreak prompt. We measure their semantic similarity using the ensemble CLIP-score [58, 92, 69], as shown in Table 5. Note that our measurement exclusively targets the examples where the defense mechanism is successful.
We calculate the CLIP-score between the original prompts $P_{query}$ from JailbreakBench and the attack-generated jailbreak prompts, as well as the CLIP-score between $P_{query}$ and our identified harmful content (prompts/intentions). Specifically, the CLIP-scores for identified harmful portions are higher than those for the generated attack prompts, suggesting that the jailbreak prompts have altered the original prompts, whereas the harmful portions identified by our models are more aligned with the original prompts. Regarding the differences between the two tuned models, we observe that the similarity of identified harmful prompts is higher than that of harmful intentions in the context of human-based, optimization-based, and generation-based attacks, but lower than that of intentions under indirect and multilingual attacks. This can be attributed to the fact that human-based, optimization-based and generation-based attack prompts contain the original prompts, whereas indirect and multilingual attack prompts significantly distort the original prompts. In summary, these findings are consistent with our initial design intentions and demonstrate that the harmful portions identified by our models can indeed provide an explainable foundation for defense mechanisms.
Table 6: Performance (ASRs) of SelfDefend with different shadow models against adaptive attacks (i.e., PAIR and TAP).
Target Model Shadow Model PAIR TAP
GPT-3.5 Llama Guard 0.38 0.39
Llama Guard 2 0.31 0.36
$P_{direct}$-tuned model 0.22 0.20
$P_{intent}$-tuned model 0.25 0.18
GPT-4 Llama Guard 0.28 0.24
Llama Guard 2 0.24 0.20
$P_{direct}$-tuned model 0.20 0.19
$P_{intent}$-tuned model 0.24 0.15
Llama-2-7b-chat Llama Guard 0.37 0.36
Llama Guard 2 0.28 0.28
$P_{direct}$-tuned model 0.21 0.21
$P_{intent}$-tuned model 0.23 0.22
@@ 7.5 Robustness to Adaptive Jailbreaks
Robustness to Black-box Jailbreaks. In §7.2, the jailbreak prompts used for evaluation are either manually designed, transferred from a surrogate model (Vicuna-7b-v1.3 [13]), or generated for the target models (GPT-3.5, GPT-4, Llama-2-7b-chat). To evaluate the robustness of our SelfDefend against black-box adaptive attacks, we select query-based PAIR and TAP to jailbreak the whole SelfDefend framework. The results are reported in Table 6. First, we can observe that all our tuned models outperform the Llama Guard series, except for one case, which is that the result of $P_{intent}$-tuned model is equal to Llama Guard 2 under PAIR. Moreover, the $P_{direct}$-tuned model is superior to the $P_{intent}$-tuned model in defending against PAIR, whether for GPT-3.5, GPT-4, or Llama-2. In contrast, the $P_{intent}$-tuned model is superior to the $P_{direct}$-tuned model in resisting TAP, except on Llama-2.
@@
Table 7: ASR with the regular GCG attack on Llama-2 and the white-box GCG directly on our tuned defense models.
Tuned Model GCG (from Table 4) White-box GCG
$P_{direct}$-tuned model 0.06 0.08
$P_{intent}$-tuned model 0.11 0.09
@@ 8 Conclusion
Ethics and the Open Science
Our research has meticulously addressed various ethical considerations to ensure responsible and ethical conduct. Firstly, regarding vulnerability disclosures and open research artifacts, we have released our datasets, raw evaluation results, and code at this anonymous GitHub link: https://github.com/SelfDefend. While our study required testing various jailbreak attacks on the live system (i.e., ChatGPT API service), these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server. We also aimed to minimize the overhead for OpenAI by distributing our experiments over a period of three months and incurred around \$2,000 in API usage fees. Secondly, in addition to all tests on ChatGPT complying with the relevant terms of service, we adhere to Meta Llama’s relevant open-source agreements to maintain legal and ethical standards. The well-being of our team members was a priority, with measures in place to protect against exposure to harmful content during the research process and offering psychological support if needed. Although our novel defense mechanism provide an effective, low-latency, plug-and-play, and explainable solution, its potential negative outcome could potentially be used maliciously to craft better attacks on LLMs. We therefore engage with the community to ensure the technology is used responsibly. This involved a readiness to retroactively identify negative outcomes and take corrective actions if our initial assessments underestimated the impacts. Finally, all aspects of our research were conducted in strict compliance with the law, ensuring that our practices did not inadvertently contravene legal standards, particularly in data protection and privacy. This comprehensive ethical approach not only aligns with the guidelines but also reinforces the integrity and societal value of our research.
@@
Appendix A Case Studies
@@
Successful Cases. Figure 5 in §7.2 presents a detection example of our methods applied to a jailbreak query, compared with the Llama Guard series. The attack prompt involves creating a fake news article to spread disinformation about a political candidate. Both Llama Guard and Llama Guard 2 identify it as safe, while our methods mark its harmful portions. The reason our method outperforms Llama Guards might be due to the enhanced reasoning ability of the foundation model in identifying harmful portions.
@@
As described in our design, the direct prompt-based model identifies harmful parts directly from the input text, whereas the intent prompt-based model summarizes the request’s intention and then identifies unsafe parts from that summary. When comparing our $P_{direct}$-tuned models with $P_{direct}$-based GPT-4, we observed that the parts they lifted from the attack prompt came from different sentences, since $P_{direct}$ emphasizes that it is sufficient to output one harmful part. A similar phenomenon occurs with the intent prompt, where the semantics of GPT-4 and our tuned model’s intercepts are not identical.
@@
In addition, please refer to Appendix C for more case of our models handling other jailbreaks or the normal prompts. As expected, both of our tuned models make satisfactory decisions on various attack or benign inputs.
@@
Appendix B Keywords
Table 8: The refusal signals considered for GPT-3.5/4.
“I’m sorry”
“Sorry”
“I am sorry”

### v3

@@ 7.3 Extra Delay $\Delta d$
Extra Delay $\Delta d$ Across Different Queries. Similar to how we measured the extra delay $\Delta d$ introduced by GPT-based SelfDefend in §5.3, we now report the average delay $\Delta d$ introduced by the tuned models on Llama-2-7b-chat in Figure 6. Compared to earlier results in Figure 3 and Figure 5, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.008 seconds. In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.072 seconds (0.026 seconds) under AlpacaEval. This is likely because the parameters of fine-tuned models are less than those of GPT-3.5/4. Besides normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under jailbreak prompts. For example, the maximum $\Delta d$ now decreases from the earlier 1.56 seconds on GPT-4 to 0.39 seconds. Indeed, except for DAN and LLM-Fuzzer, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend. These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for potential deployment.
$\Delta d$ Compared with Existing Defense Methods. We further compare the extra delay $\Delta d$ introduced by our tuned models with other defense methods. Table 6 shows the mean extra delay of different defense methods across all 11 tested jailbreak/normal queries. We observe that the extra delay $\Delta d$ of SelfDefend is significantly superior to that of other defense methods, benefiting from the parallel processing between the normal and shadow stack. Because ICD only adds in-context demonstrations to input prompts, it introduces latency closest to that of SelfDefend. Perplexity Filter and Llama Guards have much higher delays than SelfDefend, because the target models are required to wait for their jailbreak detection based on LLM inference. SafeDecoding has a much more noticeable delay because it requires two LLMs to perform simultaneous reasoning, including the target model and its fine-tuned expert model. SmoothLLM has the highest delay among all defenses, as it perturbs the input prompt and accesses the target LLM multiple times.
Table 6: The mean extra delay $\Delta d$ of different defense methods for all 11 kinds of tested jailbreak/normal inputs.
Defense Method Extra Delay $\Delta d$ (s)
$P_{direct}$-tuned SelfDefend 0.032
ICD 0.056
$P_{intent}$-tuned SelfDefend 0.077
Perplexity Filter 0.168
Llama Guard 2 0.256
Llama Guard 3 0.285
Llama Guard 0.510
SafeDecoding 0.869
SmoothLLM 21.807
@@ 7.4 Explainability
Figure 7 (see Appendix C) demonstrate the effectiveness of the tuned models in identifying harmful content within jailbreak prompts. Here, we further quantitatively assess the alignment of identified harmful content with the original jailbreak prompts by measuring semantic similarity using the ensemble CLIP-score [59, 95, 71], as shown in Table 5. Note that our measurement exclusively targets the examples where the defense mechanism is successful.
We compute the CLIP-score between the original prompts $P_{query}$ from JailbreakBench and both the attack-generated jailbreak prompts and our identified harmful content (prompts/intentions). Higher CLIP-scores for identified harmful portions compared to attack prompts indicate that while jailbreak prompts alter the original content, the harmful content identified by our models remains closely aligned with the originals.
Comparing these two tuned models, identified harmful prompts show higher similarity than harmful intentions in human-based, optimization-based, and generation-based attacks, but lower similarity in indirect and multilingual attacks. This is because the former attack types retain the original prompts, whereas the latter significantly distort them.
In summary, these findings confirm that our models effectively identify harmful content in an explainable manner, supporting robust defense mechanisms.
@@ 7.5 Robustness to Adaptive Jailbreaks
To test SelfDefend’s robustness against adaptive attacks, we consider three schemes of adaptive jailbreaks as follows.
Table 7: ASRs of SelfDefend with different shadow models against adaptive attacks (i.e., PAIR, TAP, and LLM-Fuzzer).
Target Model Shadow Model PAIR TAP LLM-Fuzzer
GPT-3.5 Llama Guard 0.38 0.39 0.61
Llama Guard 2 0.31 0.36 0.38
Llama Guard 3 0.23 0.29 0.15
$P_{direct}$-tuned model 0.22 0.20 0.17
$P_{intent}$-tuned model 0.25 0.18 0.14
GPT-4 Llama Guard 0.28 0.24 0.23
Llama Guard 2 0.24 0.20 0.07
Llama Guard 3 0.23 0.24 0.07
$P_{direct}$-tuned model 0.20 0.19 0.08
$P_{intent}$-tuned model 0.24 0.15 0.06
Llama-2-7b Llama Guard 0.37 0.36 0.10
Llama Guard 2 0.28 0.28 0.04
Llama Guard 3 0.32 0.24 0.01
$P_{direct}$-tuned model 0.21 0.21 0.00
$P_{intent}$-tuned model 0.23 0.22 0.01
Mistral-7B Llama Guard 0.43 0.41 0.37
Llama Guard 2 0.34 0.31 0.59
Llama Guard 3 0.30 0.21 0.06
$P_{direct}$-tuned model 0.27 0.24 0.10
$P_{intent}$-tuned model 0.20 0.18 0.02
Robustness to Entire Adaptive Jailbreaks. In §7.2, the jailbreak prompts used for evaluation are either manually designed, transferred from a surrogate model (Vicuna-7b-v1.3 [15]), or generated for the target models (GPT-3.5, GPT-4, Llama-2-7b-chat, and Mistral-7B-Instruct-v0.2). To evaluate the robustness of SelfDefend against black-box adaptive attacks, we select PAIR, TAP, and LLM-Fuzzer to jailbreak the entire SelfDefend. The results are reported in Table 7. First, we observe that the $P_{direct}$-tuned and $P_{intent}$-tuned models achieve the lowest ASR, outperforming the Llama Guard series. Moreover, the $P_{direct}$-tuned model is superior to the $P_{intent}$-tuned model in defending against PAIR for GPT-3.5, GPT-4, and Llama-2, except for Mistral-7B-Instruct-v0.2. In contrast, the $P_{intent}$-tuned model is superior to the $P_{direct}$-tuned model in resisting TAP and LLM-Fuzzer, except for Llama-2.
@@
We use GCG, RLbreaker, and LLM-Fuzzer to separately attack SelfDefend on Llama-2-7b-chat. For GCG, we craft the jailbreak prompt in the form “[Prefix 1] [Prefix 2] [Goal]”, where “[Prefix 1]” is optimized by GCG to attack the shadow model, and “[Prefix 2]” is optimized by GCG to attack the target model. For RLbreaker and LLM-Fuzzer, we first jailbreak the target model, then focus on attacking the shadow model with the jailbreak prompt. The results are reported in Table 8. We observe that such separate adaptive attack has a noticeable effect on the target LLM, with GCG achieving an ASR of over 40%, while RLbreaker and LLM-Fuzzer both exceed 70%. However, the ASRs of SelfDefend are significantly lower than those of the target model. The average ASR of the $P_{direct}$-tuned SelfDefend under these three jailbreaks is 0.193, while the average ASR of the $P_{intent}$-tuned SelfDefend is 0.023. This indicates that SelfDefend demonstrates encouraging robustness against separate adaptive attacks.
Table 8: ASRs (target LLM/SelfDefend) of SelfDefend with shadow models against separate adaptive attacks.
Method GCG RLbreaker LLM-Fuzzer
$P_{direct}$-tuned SelfDefend 0.43/0.32 1.00/0.15 0.76/0.11
$P_{intent}$-tuned SelfDefend 0.47/0.04 0.83/0.02 0.76/0.01
[…]
@@ 7.6 Robustness to Prompt Injection
Table 9: ASRs of individual attacks on detector models.
Detector GCG RLbreaker
Llama Guard [26] 1.00 0.74
LLM Self Defense [56] 0.99 0.82
$P_{direct}$-tuned model 0.08 0.00
$P_{intent}$-tuned model 0.09 0.00
8 Discussion
SelfDefend vs. Guardrail. As briefly introduced in §4, a guardrail approach makes direct safety judgments on the harmfulness of the content itself. It follows the typical classification mindset to determine whether the given input (typically LLMs’ responses) is harmful or not, without considering the target models and utilizing their inherent safety alignment. SelfDefend, on the other hand, designs the defense from the perspective of a target model, using one instance of the target model to process the input as usual and using another instance (which could be the target model itself, as we demonstrated in §5, or a dedicated defense model described in §6) to activate the detection state simultaneously. Since the detection state in the shadow stack could collaborate with the answering state in the normal stack and still utilize its safety alignment, SelfDefend’s approach would theoretically have higher defense effectiveness than any guardrail methods, as we empirically tested in §7.2, while also have minimal extra delay, as shown in §7.3. Therefore, SelfDefend’s architecture should be preferred over the guardrail approach.
Moreover, SelfDefend’s tuned models consider direct malicious portions or indirect malicious intentions first before making safety judgments, which better captures the essence of jailbreak detection since jailbreaking must involve malicious goals. Thus, our shadow models are more effective against jailbreak prompts (see §7.2) and less likely to be hacked (see §8) compared to the Llama Guard series, which make safety judgments first and then explain them—from effect to cause.
Impact on Utility. We clarify that the utility of SelfDefend is not compromised by its defense effectiveness. First, SelfDefend is a plug-in, self-contained defense method that does not require any modifications to the target model. It is thus easy to deploy and does not require altering the output of the target LLM. Second, as previously illustrated in Table 3 and Table 4, the shadow model-enabled framework shows similar ASRs to the target LLM on the normal prompt dataset AlpacaEval. Since AlpacaEval contains about 85 instructions on code generation and 720 samples on daily question-answering, it is expected that the normal user experience would not be affected by SelfDefend’s defense.
@@ 9 Conclusion
Acknowledgements
We thank the reviewers and the shepherd for their constructive comments. The HKUST authors are supported in part by an RGC CRF grant under contract C6015-23G, a research fund provided by HSBC, and a UGC TLIP project (175X-2425). The NTU authors are supported by the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2GC-2023-008), and NRF Investigatorship NRF-NRFI06-20200001. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Cyber Security Agency of Singapore.
Ethics Considerations
Our research has meticulously addressed various ethical considerations to ensure responsible and ethical conduct. Firstly, while our study required testing various jailbreak attacks on the live system (i.e., ChatGPT API service), these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server. We also aimed to minimize the overhead for OpenAI by distributing our experiments over a period of three months and incurred around \$2,000 in API usage fees. Secondly, in addition to all tests on ChatGPT complying with the relevant terms of service, we adhere to Meta Llama’s relevant open-source agreements to maintain legal and ethical standards. The well-being of our team members was a priority, with measures in place to protect against exposure to harmful content during the research process and offering psychological support if needed. Although our novel defense mechanism provides an effective, low-latency, plug-and-play, and explainable solution, its potential negative outcome could potentially be used maliciously to craft better attacks on LLMs. We therefore engage with the community to ensure the technology is used responsibly. This involved a readiness to retroactively identify negative outcomes and take corrective actions if our initial assessments underestimated the impacts. Finally, all aspects of our research were conducted in strict compliance with the law, ensuring that our practices did not inadvertently contravene legal standards, particularly in data protection and privacy. This comprehensive ethical approach not only aligns with the guidelines but also reinforces the integrity and societal value of our research.
Open Science
We have released our datasets, raw evaluation results, and code at this GitHub link: https://github.com/SelfDefend. We also release the artifact at https://doi.org/10.5281/zenodo.14736935.
@@ Acknowledgements
Appendix A Expanded Evaluation across Other LLMs
As shown in Table 10, we extend the evaluation of our tuned models to other LLMs, including Claude-3.5-sonnet and Llama-2-13b-chat. For the Claude-3.5-sonnet API, we use the version “claude-3-5-sonnet-20241022”. As in the earlier evaluations on GPT-3.5/4, we use a similar setting to generate jailbreaks for Claude. For Llama-2-13b-chat, we use an analogous setting to Llama-2-7b-chat in prior evaluations. Under the target LLM Claude-3.5-sonnet, we observe that all jailbreak methods have low ASRs, indicating the resistance of this newest Claude. Despite this, SelfDefend, either $P_{direct}$-tuned or $P_{intent}$-tuned, achieves the lowest ASR among all defense methods, demonstrating the effectiveness of our tuned models. On Llama-2-13b-chat, our methods also achieve the lowest ASRs, except for ICD under PAIR. For normal prompts, SelfDefend shows similar ASRs to the target LLMs, indicating that the utility of the target LLMs is not compromised by our defense.
Table 10: Jailbreak ASRs for various defense methods. For ICD and SafeDecoding, we present the performance of their enhanced models. For detection-based Perplexity Filter, SmoothLLM and Llama Guards, we report ASRs only on their detection modules.
Model DAN GCG AutoDAN RLbreaker PAIR TAP LLM-Fuzzer DrAttack Puzzler MultiJail AlpacaEval
Claude Claude-3.5-sonnet (baseline) 0.029 0.020 0.010 0.300 0.280 0.260 0.110 0.160 0.010 0.048 0.971
ICD [75] 0.095 0.010 0.000 0.060 0.160 0.160 0.040 0.140 0.120 0.333 0.903
Perplexity Filter [27] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [60] 0.103 0.030 0.000 0.080 0.160 0.190 0.020 0.160 0.030 0.305 0.958
Llama Guard [26] 0.552 0.400 0.560 0.610 0.480 0.420 0.690 0.980 0.930 0.952 0.996
Llama Guard 2 [67] 0.432 0.140 0.200 0.370 0.380 0.360 0.240 0.910 0.620 0.559 0.990
Llama Guard 3 [19] 0.343 0.040 0.070 0.020 0.270 0.270 0.010 0.620 0.420 0.378 0.986
(3.5-sonnet) $P_{direct}$-tuned Shadow Stack 0.260 0.060 0.060 0.020 0.240 0.290 0.020 0.740 0.100 0.737 0.960
$P_{direct}$-tuned SelfDefend 0.025 0.010 0.000 0.010 0.170 0.160 0.000 0.160 0.000 0.044 0.939
$P_{intent}$-tuned Shadow Stack 0.296 0.070 0.060 0.040 0.190 0.210 0.030 0.170 0.150 0.613 0.993
$P_{intent}$-tuned SelfDefend 0.025 0.020 0.000 0.020 0.150 0.140 0.020 0.080 0.000 0.038 0.966
Llama-2 Llama-2-13b-chat (baseline) 0.761 0.640 0.780 0.610 0.320 0.280 0.160 0.670 1.000 0.181 0.984
ICD [75] 0.593 0.390 0.680 0.500 0.140 0.150 0.620 0.340 0.980 0.270 0.845
SafeDecoding [84] 0.753 0.900 0.850 0.650 0.300 0.300 0.660 0.700 1.000 1.000 0.983
Perplexity Filter [27] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [60] 0.911 0.850 0.950 0.920 0.600 0.590 0.840 0.690 1.000 0.321 0.996
Llama Guard [26] 0.552 0.450 0.950 0.730 0.430 0.340 0.540 0.890 0.930 0.952 0.996
Llama Guard 2 [67] 0.432 0.230 0.100 0.280 0.390 0.360 0.250 0.890 0.620 0.559 0.990
Llama Guard 3 [19] 0.343 0.110 0.080 0.110 0.290 0.210 0.150 0.450 0.420 0.378 0.986
(13b-chat) $P_{direct}$-tuned Shadow Stack 0.256 0.120 0.060 0.020 0.270 0.170 0.020 0.400 0.130 0.740 0.963
$P_{direct}$-tuned SelfDefend 0.221 0.120 0.060 0.000 0.250 0.130 0.000 0.270 0.130 0.121 0.949
$P_{intent}$-tuned Shadow Stack 0.286 0.090 0.050 0.060 0.270 0.190 0.050 0.120 0.200 0.619 0.990
$P_{intent}$-tuned SelfDefend 0.244 0.090 0.050 0.030 0.220 0.150 0.010 0.100 0.200 0.111 0.975
Appendix B Average $\Delta d$ for Various Jailbreak Prompts
Figure 5: Average $\Delta d$ introduced by GPT-based SelfDefend across various jailbreaks.
Figure 6: Average $\Delta d$ introduced by SelfDefend’s tuned models on Llama-2-7b-chat.
Appendix C Case Studies
@@
Successful Cases. Figure 7 presents a detection example of our methods applied to a jailbreak query, compared with the Llama Guard series. The attack prompt involves creating a fake news article to spread disinformation about a political candidate. Both Llama Guard and Llama Guard 2 identify it as safe, while our methods mark its harmful portions. The reason our method outperforms Llama Guards might be due to the enhanced reasoning ability of the foundation model in identifying harmful portions.
@@
As described in our design, $P_{direct}$-based model identifies harmful parts directly from the input text, whereas $P_{intent}$-based model summarizes the request’s intention and then identifies unsafe parts from that summary. When comparing our $P_{direct}$-tuned models with $P_{direct}$-based GPT-4, we observed that the parts they lifted from the attack prompt came from different sentences, since $P_{direct}$ emphasizes that it is sufficient to output one harmful part. A similar phenomenon occurs with the intent prompt, where the semantics of GPT-4 and our tuned model’s intercepts are not identical.
@@
In addition, please refer to https://github.com/selfdefend/DatasetAndResult/blob/main/examples.md for more case of our models handling other jailbreaks or normal prompts. As expected, both of our tuned models make satisfactory decisions on various attack or benign inputs.
@@
Appendix D Keywords
Table 11: The refusal signals considered for GPT-3.5/4 and Claude.
“I’m sorry”, “Sorry”, “I am sorry”